OpenStudy (anonymous):

In a certain right triangle, the one leg is two times longer than the other leg and the the hypoteneuse is four feet longer than the shorter leg. If x represents the length of the shorter leg, then by the Pythagorean Theorem this problem translates to which equation below: a. x2 + 2x2 =4x2 b. x2 + (x + 4)2 =2x2 c. x2 + (x + 4)2 =4x2 d. x2 + 2x2 =(x + 4)2 e. x2 + 4x2 =(x + 4)2

OpenStudy (sriram):

E

OpenStudy (anonymous):

could you explain how you get that?

OpenStudy (sriram):

the two legs which aren't hypotnuseare x and 2x hypotnuse=x+4 hencex^2+(2x)^2=(x+4)^2 giving x^2+4x^2=(x+4)^2

OpenStudy (anonymous):